You are a manager at the Donnelly Corporation, a mirror and window supplier to the major automakers. Recently, you conducted a study of the production process for your DirectBond single-side encapsulated window (a product that was first introduced on Chrysler minivans). The results from the study are summarized in the table below, and are based on the 5 units of capital currently available at your plant. Workers are paid $50 per unit, per unit capital costs are $10, and your encapsulated windows sell for $5 each. Given this information, optimize your human resource and production decisions. Do you anticipate earning a profit or a loss? Explain carefully.

labour output

0 0

1 10

2 30

3 60

4 80

5 90

6 95

7 95

8 90

9 80

10 60

11 30


1
Expert's answer
2021-04-08T07:26:20-0400

"\bigstar"

"\bull" The mirror and window supplier is a profit-maximizing firm that will hire labor up to the point where the value of its marginal product equals the wage rate of $50.


"\bull" There'll are currently 5 units of capital available at the plant.


"\bull" Firm’s demand for labor is the only decision that the firm is conscious about. Given the wage rate of $60 per unit of labor,


"\bull" the firm will continue to hire labor till the value of marginal product of labor reaches $60.


labour output MP AP


0 0 0 0


1 10. 10-0=10 10/1=10


2 30. 30-10=20. 30/2=15


3 60. 60-30=30. 60/3=20


4 80. 80-60=20. 80/4=20


5 90. 90-80=10. 90/5=14


6 95. 95-90=5. 95/6=15,8


7 95. 95-95=0. 95/7=13,5


8 90. 90-95=-5. 90/8=11,25


9 80. 80-90=-10. 80/9=8,89


10 60. 60-80=-20. 60/10=6


11 30. 30-60=-30. 30/11=2,9



"\bigstar" Thus maximizing productivity and return on resources Observed when hiring 3 and 4 employees Profit maximization will be observed if 4 employees are hired:



80 * 5-50 * 4-20 * 5 = $ 100